A Space & astronomy forum. SpaceBanter.com

Go Back   Home » SpaceBanter.com forum » Space Science » Space Station
Site Map Home Authors List Search Today's Posts Mark Forums Read Web Partners

ISS On-Orbit Status, 17-03-2004



 
 
Thread Tools Display Modes
Prev Previous Post   Next Post Next
  #1  
Old March 19th 04, 08:57 PM
Jacques van Oene
external usenet poster
 
Posts: n/a
Default ISS On-Orbit Status, 17-03-2004

ISS On-Orbit Status 17 Mar 2004

All ISS systems continue to function nominally, except as noted previously
or below.

Early in the morning, FE Alexander Kaleri concluded Part 3 of his current
MBI-8 "Profilaktika" (countermeasures) fitness assessment, first, in the
morning, with the blood tests (to determine lactate and creatine kinase
levels in the blood with the AccuSport equipment), then, later in the day,
by the physical exercise session on the TVIS treadmill. CDR Foale was
available to assist as CMO (crew medical officer) as required. [The TVIS
test is identical to the MO-3 test performed on the treadmill in idling mode
with free choice of speeds within a certain specified range. In addition to
the nominal test procedure, MBI-8/Part 3 calls for the use of the TEEM-100M
gas analyzer during the test, the blood lactate measurements, and subjective
evaluation of physical exertion levels during the test. At the end of the
creatine kinase tests, the results were logged, copied from
Cardiocassette-2000 recording to OCA for downlink, and reported to the
ground.]

FE Kaleri then made preparations for tomorrow's U.S. IMG PHS (Integrated
Medical Group/Periodic Health Status) tests without blood labs, which is
part of the comprehensive medical evaluation, by setting up the required
equipment near the medical equipment computer (MEC). [PHS w/o Blood Labs
tests are performed with equipment from the crew's ambulatory medical pack
(AMP) and oral disposable thermometers, delivered on Progress 13P as part of
the IMAK (ISS medical accessory kit), guided by special software (IFEP,
in-flight examination program) on the MEC. All PHS data are to be recorded
on the IFEP, using the same forms used for MO-9.]

Foale completed the mandatory 30-min. CBT (computer-based training with
video and audio) to refresh his Crew Medical Officer (CMO) proficiency,
deferred from 3/15. [To maintain proficiency in using HMS (health
maintenance systems) hardware including ACLS (advanced cardiac life support)
in contingency situations where crew life is at risk, these training
sessions are performed once a month for one hour to review equipment and
procedures via CBT (computer-based training). Besides ACLS, procedures
include airway obstruction management, i.e., review of suction device, nasal
airway, intubating laryngeal mask airway (ILMA) with endotracheal tube, and
cricothyrotomy (incision to re-enable breathing air inflow).]

With new SVO-ZV hose adapters delivered on 13P, Michael Foale undertook the
nominally monthly (Week 21) potable water sampling for in-flight
chemistry/microbiology analysis, using approved Russian sampling procedures
with the U.S. WS&A (water sampler & archiver) for collection and the WMK
(water microbiology kit) for treatment within 6 hours of the collection.
Results will be available after a two-day incubation period. [Samples were
taken in the SM at the potable water SRV-K hot port and from the EDV
container of the SVO-ZV water supply system, using new media syringes sent
up on 13P. Last time done: 12/3/03.]

After the HRF (human research facility) rack was powered up by the ground,
Foale transferred the FOOT (Foot/Ground Reaction Forces during Space Flight)
experiment data of his most recent session, Mike's third (3/12), onto the
HRF PC (portable computer), created content files for ground analysis, and
set up the laptop for downlink operations.

Also on the HRF, the CDR checked out the MedOps cardiac defibrillator, a
periodic routine task that is scheduled as soon as possible from Expedition
start and every 60 days thereafter. [For the checkout, the defib is
connected to the 120V outlet, equipped with its battery, today #1007, and
then allowed to charge, for about five seconds, to a preset energy level
(e.g., 100 joules). After the button-triggered discharge, a console
indicator signals success or failure of the test. The pacing signal was to
be downlinked via S-band for 2 min. The HRF was powered down after the
checkout.]

Kaleri worked on the TP1 laptop in the SM, installing a new software
application (named FHS View, from Rohde & Schwarz/Germany) for the
Russian/German Global Timing System (GTS). [GTS electronics are installed
in the SM, with two transmitters broadcasting a time signal at 400.1 MHz (1
watt) and 1428 MHz (0.5 watt) straight downward (nadir) for worldwide clock
synchronization. The system has apparently been experiencing some problems
for a while.] Sasha performed his regular maintenance on the BIO-5
Rasteniya-2 ("Plants-2") payload, "refueling" its water storage tank as
required. [Rasteniya studies growth and development of plants (peas) under
spaceflight conditions in the Lada-4 greenhouse. Regular maintenance
involves monitoring of seedling growth, humidity measurements, watering to
moisten the substrate if necessary, and photo/video recording.]

In preparation for tomorrow's robotics operations with the SSRMS (Space
Station Remote Manipulator System), Mike connected the UOP-DCP (utility
outlet panel-to-display & control panel) bypass power cable at the Lab RWS
(robotics workstation). The ground then powered up the Lab RWS, the SSRMS
on the redundant string and the VDUs (video distribution units), but not the
cameras as yet. [In addition, via ground commanding, an UMA (umbilical
mating assembly) data downlink from the primary EXT MDM (external
multiplexer/demultiplexer, computer) was executed. More data "dumps" are
scheduled for tomorrow during the crew training.]

The FE conducted the periodic status checkup of the airflow sensors (IP-1),
which are installed in the various RS (Russian segment) hatchways and the
FGB-to-Node hatchway. [Last time done: 11/12/03.]

Alex conducted the routine technical maintenance of the SM's SOZh
environment control & life support system, incl. the toilet equipment (ASU),
and prepared the daily IMS "delta" file for automatic export/import to
update the database.

Sasha also conducted an inventory audit in the DC-1 docking compartment,
checking on the contents of bags to find equipment listed as "missing".
Primary attention is on 12 EVA gloves; if necessary, the search was to be
extended to the Service Module (SM) and FGB. [The FE was authorized to
reconfigure structural elements (panels) in the DC-1 and use its handrails
as required to facilitate equipment relocation.]

The Elektron water electrolysis system continues to be inoperative.
TsUP/Moscow, after discussions with MCC/Houston, has now decided to replace
the Elektron's Liquid Unit (BZh-6) with the remaining on-board spare. The
R&R will be scheduled ASAP. [Two theories were advanced for the cause of
the ketone-like odor reported by the crew: (a) electrolyte (KOH, potassium
hydroxide) coming in contact with rubber in a water pump used temporarily to
add water to the BZh, which could have created a new compound that was then
processed by the Elektron and produced the smell; (b) hydrogen (H2) seepage
into the catalyst/charcoal bed purification unit resulting in a
hotter-than-usual charcoal filter, which then may have released bad-smelling
contaminants.]

To raise ppO2 (oxygen partial pressure) in the cabin atmosphere, two more
SFOG candles were burned today. This makes a total of eight candles used
since last Saturday. None of the candles experienced any anomalies. [The
SFOGs (solid-fuel oxygen generators), on standby as oxygen sources as backup
to the Elektron and to Progress-stored O2, generate oxygen by decomposing
cartridges of solid lithium perchlorate (LiClO4) into lithium chloride
(LiCl) and O2 when heated at 400 degC. Each candle produces 600 liters
(1.74 lb) of O2, enough for one person/day.]

At 9:46 am EST, the crew supported a 20-min. interactive PAO/educational TV
event with Howard Bishop Middle School in Gainesville, FL. [The school is
participating in the inaugural year of the NASA Explorers Schools program
and intends to systematically extend NASA content from, initially, the
school of 1200 middle school students to the surrounding elementary students
and community members. During today's event, 400 participants were in the
main auditorium of the school, and the program was broadcast throughout the
entire school as well as distributed on the local Cox Cable TV network.]

At 9:28am, Mike conducted a ham radio session with students at Bhac School
on the Isle of Lewis, part of Scotland's Western Isles. [In the primary
department, students are taught in Gaelic or in English language, while in
the secondary department all classes are in the English.]

Instructions were uplinked for recovering comm between the MSG (Microgravity
Science Glovebox) and its MLC (MSG Laptop Computer) by verifying proper
installation of cards and cable connections.

Onboard consumables/resources continue to be monitored carefully. Water
supplies are tracking just above the (adjusted) planning line. Oxygen is
well below predicted consumption since Elektron remains down; future O2 plan
is currently being evaluated.

Early this morning (3:27am EST), CMG-3 vibrations showed a very brief spike
exceeding the 0.06g limit. CMG performance continues to be studied closely.

SM storage battery #8 is off-line, leaving seven of the 800A accumulators
functioning nominally. [Battery #8 began discharging on its own last
night, about 90 minutes before a planned battery cycling, and did not
respond to program commands. Its current converter control unit (BUPT-1M)
is past its service life and may be the cause. A space BUPT is on board.]

Follow-up to yesterday's report on Caution & Warning alarms during the fire
emergency drill: C&W experts have now determined that the commands to
silence the alarm worked properly. No anomaly will be documented for this
case.

Today's CEO (Crew Earth Observations) targets, limited in XPOP attitude by
flight rule constraints on the use of the science window, which is available
for only ~1/4 of each orbit when not facing forward (in "ram"), were Perth,
Australia (nadir pass. Perth lies inland from its port of Freemantle on the
Swan River. With 1.3 million people, the Perth-Freemantle urban zone is the
center of farming and business activity in Australia's west), Cape Town,
South Africa (nadir pass. The city sprawls tens of kilometers eastward onto
flat sandy plains [between the peninsula mountains, where Cape Town is
centered, and the main mountain ranges inland]), Johannesburg, South Africa
(nadir pass. Suggested views were just left and just right of track),
Internal waves, Falkland (views just left may reveal internal waves near the
glint point. Shallow sea floor stretches 3000 kilometers east from
southernmost South America into the Atlantic, with the Falklands and South
Georgia archipelago as high points. This entire region thus has research
interest for internal waves), and Lima, Peru (nadir pass over another city
that lies inland from its port [Callao]).

CEO images can be viewed at the websites.

http://eol.jsc.nasa.gov
http://earthobservatory.nasa.gov

See also the website "Space Station Challenge" at
http://voyager.cet.edu/iss/

U.S. and Russian Segment Status (as of today, 1:49pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

Elektron O2 generator is Off. Vozdukh CO2 scrubber is On. U.S. CDRA CO2
scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant
control subsystem) is operating. SM Gas Analyzer has been calibrated and is
used for ppO2 and ppCO2 monitoring. MCA (major constituents analyzer) is in
Life Extending Mode (LEM). BMP Harmful Impurities unit: absorbent bed #1 in
Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On, SKV-2
is Off (repair now completed; to be tested ASAP).


SM Working Compartment: Pressure (mmHg) -- 742; temperature (deg C) --
26.4; ppO2 (mmHg) -- 140.0; ppCO2 (mmHg) -- 3.7;
SM Transfer Compartment: Pressure (mmHg) -- 743; temperature (deg C) --
20.8.
FGB Cabin: Pressure (mmHg) -- 744; temperature (deg C) -- 24.0.
Node: Pressure (mmHg) -- 740.7; temperature (deg C) -- 24.4 (shell); ppO2
(mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
U.S. Lab: Pressure (mmHg) -- 742.49; temperature (deg C) -- 24.7; ppO2
(mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
Joint Airlock (Equip. Lock): Pressure (mmHg) -- 742.40; temperature (deg
C) -- 28.5; shell heater temp (deg C) -- 26.9, ppO2 (mmHg) -- n/a; ppCO2
(mmHg) -- n/a.
PMA-1: Shell heater temp (deg C) -- 24.0
PMA-2: Shell heater temp (deg C) -- 9.8

(n/a = data not available)

Electrical Power Systems (EPS):

Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B
both in Autotrack (suntracking) and bias-angled 43 deg. for drag reduction
("sun slicer")
SM batteries: Battery #8 is off-line; all other batteries (7) are in
"Partial Charge" mode.
FGB batteries: Battery #6 is off-line (capacity restoration mode, ROM);
battery #4 is in "Cycle" mode; all other batteries (4) are in "Partial
Charge" mode.
Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
INT-2 is operating; INT-1 is Off.
EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is Off; PL-2 MDM is Operational.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1
dropped out 11/22/03).
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3
dropped out 10/22/03).

Propulsion System:

Total propellant load available: 3948 kg (8704 lb) as of 3/5/04 [SM(755) +
FGB(2534) + Progress M(0) + Progress M-1(659)]. (Capability: SM -- 860 kg;
FGB -- 6120 kg).

Attitude Control Systems:

3 CMGs on-line (CMG-1 failed).
State vector source -- U.S. SIGI-1 (GPS)
Attitude source -- U.S. SIGI-1 (GPS)
Angular rate source -- RGA-1

Flight Attitude:

XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw: 0.5 deg,
pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist).

Communications & Tracking Systems:

FGB MDM-1 is powered Off; FGB MDM-2 is operational.
All other Russian communications & tracking systems are nominal.
*S-band is operating nominally (on string 2).
Ku-band is operating nominally.
Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
Video subsystem operating nominally.
HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

SSRMS/Canadarm2 based at Lab PDGF/LEE A, powered on both strings.
MBS: KA power on both strings.
MT: latched and mated at WS4.
POA: KA power on both strings.
RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is
Off.

ISS Orbit (as of this morning, 6:25am EST [= epoch]):

Mean altitude -- 367.8 km
Apogee -- 374.9km
Perigee -- 360.6 km
Period -- 91.90 min.
Inclination (to Equator) -- 51.6282 deg
Eccentricity -- 0.0010627
Solar Beta Angle -- 49 deg (magnitude decreasing)
Orbits per 24-hr. day -- 15.67
Mean altitude loss last 24 hours -- 110 m
Revolutions since FGB/Zarya launch (Nov. '98) -- 30392

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times,
see http://www.hq.nasa.gov/osf/station/viewing/issvis.html



--
---------------------------

Jacques :-)

www.spacepatches.info




 




Thread Tools
Display Modes

Posting Rules
You may not post new threads
You may not post replies
You may not post attachments
You may not edit your posts

vB code is On
Smilies are On
[IMG] code is On
HTML code is Off
Forum Jump

Similar Threads
Thread Thread Starter Forum Replies Last Post
ISS Status Report No. 06 - 2004 Jacques van Oene Space Station 0 January 29th 04 05:50 PM
ISS Status Report No. 04 - 2004 Jacques van Oene Space Station 0 January 17th 04 01:01 PM
ISS Status Report No.3 - 2004 Jacques van Oene Space Station 0 January 13th 04 10:22 AM
ISS Status Report No. 01 - 2004 Jacques van Oene Space Station 0 January 4th 04 02:43 PM
Ed Lu Letter from Space #6 Jacques van Oene Space Station 0 July 4th 03 11:10 AM


All times are GMT +1. The time now is 10:29 PM.


Powered by vBulletin® Version 3.6.4
Copyright ©2000 - 2025, Jelsoft Enterprises Ltd.
Copyright ©2004-2025 SpaceBanter.com.
The comments are property of their posters.